# Protect Measurement-Induced Phase Transition from Noise

### v2

@@ Abstract
Measurement-induced phase transition (MIPT) is a novel non-equilibrium phase transition characterized by entanglement entropy. The scrambling dynamics induced by random unitary gates can protect information from low-rate measurements. However, common decoherence noises, such as dephasing, are detrimental to the volume law phase, posing a significant challenge for observing MIPT in current noisy intermediate-scale quantum devices. Here, we demonstrate that incorporating quantum-enhanced operations can effectively protect MIPT from environmental noise. The conditional entanglement entropy is associated with a statistical mechanics model wherein noise and quantum-enhanced operations act as two competing external random fields. Then we show that an average apparatus-environment exchange symmetry ensures the conditional entanglement entropy is a valid probe of entanglement. Furthermore, we provide numerical evidence on a (2+1)-d quantum circuit under dephasing noise, demonstrating that MIPT can indeed be observed with the aid of quantum-enhanced operations. This result not only serves as a concrete example of the power of quantum enhancement in combating noise but also holds experimental relevance, as the protocol is straightforward to implement in practice.
@@
We first provide analytical analysis, mapping the random quantum circuit to a statistical-mechanics model. We find that noise and QE operation exactly map to two symmetry-breaking fields in different directions. Combining with unitary gate and measurement, we obtain a random field model with ferromagnetic coupling that supports a ferromagnetic-paramagnetic phase transition with both increasing temperature and increasing random field strength. Next, we discover that an average apparatus-environment exchange (aAEE) symmetry ensures the conditional entanglement entropy (CEE) to be a valid probe of entanglement [43, 44]. By imposing the aAEE symmetry, we conduct numerical studies for (2+1)-d circuits and explicitly demonstrate the MIPT in the presence of various decoherence noises. Our approach is inspired by Ref. [43], where a prototype setting was introduced. The structure we introduced here incorporates measurement and is much easier to implement in practice, as will be further elaborated later.
@@
Analytical analysis. Instead of considering subsystem entanglement entropy, we aim to calculate the CEE $\overline{S(M|A)}\equiv\overline{S(M,A)}-\overline{S(A)}$. The overline here represents averaging all circuits and measurement outcomes. Mathematically, this can be expressed as,
@@
$\mathcal{F}^{n,k}=-\text{log}\mathcal{Z}^{(n,k)}$ is identified as the free energy [46]. The underlying statistical-mechanics model for $\mathcal{Z}^{(n,k)}$ is composed of spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$. We designate the identity operation in $\mathbb{S}(Q)$ as $\mathbb{I}_{Q}$ and a particular permutation operator as $\mathbb{C}=\mathbb{S}^{\otimes k}\otimes\mathbb{I}$. $\mathbb{S}$ is the cyclic permutation operation in permutation group $\mathbb{S}(n)$. Thus, $\overline{S(M|A)}$ is related to the free energy difference under different boundary conditions, as shown in Fig. 2(a). In both $\mathcal{Z}_{MA}$ and $\mathcal{Z}_{A}$, the environment qubits are always subject to the boundary condition $\mathbb{I}$ while the ancilla qubits are subject to $\mathbb{C}$. It has been shown that the entanglement phase transition can be seen as the spontaneous breaking of the $(\mathbb{S}_{Q}\times\mathbb{S}_{Q})\rtimes\mathbb{Z}_{2}$ symmetry [47, 48, 49]. However, boundary condition imposing on environment qubits can be effectively understood as an explicit symmetry-breaking field in the bulk that align the spins to direction $\mathbb{I}$, which is thus detrimental to MIPT [35, 36, 37]. Through QE operation and considering the conditional entropy, we effectively introduced another symmetry-breaking field in direction $\mathbb{C}$ to counteract with the effect of noises. Taking the example of the resetting noise and its associated symmetric QE operation $U_{SE}=U_{SA}=\text{SWAP},\rho_{E}=\rho_{A}=\left|0\right\rangle\left\langle 0\right|$, the bond weight connecting two spins $\sigma$ and $\tau$ can be explicitly formulated as [46],
$W^{\text{reset}}(\sigma,\tau)$ $=$ $(1-p)(1-q)\left\langle\sigma|\tau\right\rangle$ (3)
$+(1-p)\frac{q}{2}\left(\left\langle\sigma|\mathbb{C}\right\rangle+\left\langle\sigma|\mathbb{I}\right\rangle\right)+pd.$
We choose $q_{n}=q_{e}=q/2$ which ensures that the total symmetry-breaking field is zero. The model has a $(\mathcal{C}_{\mathbb{S}_{Q}}(\mathbb{C})\rtimes\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$ symmetry where $\mathcal{C}_{\mathbb{S_{Q}}}(\mathbb{C})$ is the centralizer of $\mathbb{C}$ [46]. Note that $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$ where $d$ is the local Hilbert space dimension and $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$. In the $d\to\infty$ limit, $\left\langle\sigma|\tau\right\rangle$ becomes $d^{Q}\delta_{\sigma,\tau}$. This corresponds to a random field $Q!$-state Potts model, where the first term can be interpreted as a ferromagnetic coupling, the second term as two competing random fields, and the final term as the effect of temperature. The strength of the random fields is controlled by $q$ while temperature is controlled by $p$. Given prior work demonstrating phase transitions in random field Ising models and Potts models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59], we anticipate that a phase transition would occur at finite $p$ and $q$ in quantum circuit which is directly characterized by $\overline{S(M|A)}$. For depolarizing and dephasing noises, the corresponding statistical-mechanics models are analogous as provided in Supplemental Material [46].
@@
Average symmetry. Since $\overline{S(M|A)}$, as opposed to $\overline{S(M)}$ in traditional setting, undergoes a phase transition, it is pertinent to determine under what conditions $\overline{S(M|A)}$ serves as a valid probe of entanglement or quantum information [60]. It has been proven that if the system has information exchange (IE) symmetry in every realization, such that $S(M|A)=S(M|E)$, then $S(M|A)$ ensures entanglement between $M$ and $M^{c}\cup A$ [44]. Here, $M^{c}$ denotes the complement of the subsystem $M$. Volume law scaling of $S(M|A)$ would then yield an extensive number of distillable Bell pairs between $M$ and $M^{c}\cup A$. The IE symmetry can be guaranteed if an apparatus-environment exchange (AEE) symmetry is present, which requires that each circuit realization be symmetric under the exchange of ancilla and environment qubits. However, this symmetry is explicitly violated by any realization in our circuit since noise and QE operation are independently chosen and applied at each location. Additionally, the number of ancilla and environment qubits may differ within a specific circuit realization, making the “exchange” operation more subtle to define. To solve this issue, we propose that an aAEE symmetry can still lead to an average IE (aIE) symmetry, thereby making the averaged CEE a valid probe of entanglement. The relationship between these symmetries are summarized in Fig. 3. We first define a generalized SWAP operation $\text{SWAP}^{\prime}$ on $\left|\Psi_{m}\right\rangle$ as initially adding unentangled qubits to either $E$ or $A$ to equalize the number of qubits. Subsequently, a conventional SWAP gate is applied to $E$ and $A$, followed by the removal of the unentangled qubits. $\text{SWAP}^{\prime}\left|\Psi_{m}\right\rangle$ would not necessarily equal to $\left|\Psi_{m}\right\rangle$, but may transform it into another realization $\left|\Psi_{m^{\prime}}\right\rangle$. The aAEE symmetry requires that the probabilities of these two realizations occurring are identical. Concretely, one can represent the circuit’s outcome as $\rho=\sum_{m}\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|$, and the aAEE symmetry necessitates that
@@
This is reminiscent of the recently proposed average (weak) symmetry in the study of mixed state order [61, 62, 63, 64, 65, 66]. A subtlety in defining $\rho$ in this manner is that the number of qubits in $\left|\Psi_{m}\right\rangle$ may vary, but this can be simply addressed by adding unentangled qubits to states with fewer qubits. With aAEE symmetry and combining with Eq. (2), one would have
$\sum_{m}\Lambda_{A}^{Q}$ $=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\right.$ (5)
$\left.\text{SWAP}^{\prime\dagger}\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It is then straightforward to show that aIE symmetry results in,
$\overline{S(M|A)}\geq 0,\quad\overline{S(M|A)}=\overline{S(M^{c}|A)},$
$\overline{S(M|A)}=-\overline{S(M|A\cup M^{c})},$ (6)
which also ensure that $\overline{S(M|A)}$ represents the entanglement between $M$ and $M^{c}\cup A$.
Translating the aAEE symmetry into the concrete circuit model reveals that it requires $\rho_{E}=\rho_{A}$, $U_{SE}=U_{SA}$ and $q_{n}=q_{e}=q/2$. Thus, QE operation should be symmetric with the corresponding noise, and the operation should be applied at the same rate as the noise. It is noteworthy that aAEE symmetry condition $q_{n}=q_{e}$ aligns with the condition of zero net field in the statistical-mechanics model. Since every noise can be represented in the Stinespring form [45], QE operations can protect MIPT from various types of noise.
@@
An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting. Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$. In the thermodynamic limit, curves corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$. A more detailed explanation and finite size numerical results are provided in [46].
Besides noise, post-selection is another significant challenge preventing MIPT from being easily observed. Here the post-selection problem persists when the measurement rate is non-zero because determining the CEE of a single trajectory requires post-selecting on the measurement outcomes, a difficulty rooted in Born’s rule. Recent work has utilized various post-selection-free methods to probe the phase transition, such as using cross-entropy benchmarking [72] or providing bounds on the entanglement entropy [73, 74]. We believe these methods can also be extended to the conditional entropy in this work, and we leave this exploration for future work.
@@ Acknowledgements.
Acknowledgment. We thank Shane P. Kelly for valuable discussions. This work is supported by the National Key Research Program of China under Grant No. 2019YFA0308404, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600 and No. 2019SHZDZX01.

### v3

@@ Abstract
Scrambling dynamics induced by random unitary gates can protect information from low-rate measurements, which underpins the phenomenon known as the measurement-induced phase transition (MIPT). However, typical decoherence noises disrupts the volume law phase, complicating the observation of MIPT on noisy intermediate-scale quantum devices. Here, we demonstrate that incorporating quantum-enhanced operations can effectively protect MIPT from environmental noise, thereby enabling its detection in experiment. The transition is characterized by the conditional entanglement entropy (CEE), which is associated with a statistical mechanics model wherein noise and quantum-enhanced operations act as competing external random fields. When the net external field is zero, a ferromagnetic-paramagnetic phase transition is expected, resulting in the MIPT. This zero-field condition also ensures an average apparatus-environment symmetry, making CEE a valid probe of entanglement and establishing the transition as a genuine entanglement phase transition. Additionally, we provide numerical results demonstrate the MIPT in a (2+1)-dimensional quantum circuit under dephasing noise. We also propose a method to estimate the noise rate, enabling the zero-field condition to be achieved experimentally and ensuring the feasibility of our protocol. Our result serves as a concrete example of the power of quantum enhancement in combating noise.
@@
We first provide analytical analysis, mapping the conditional entanglement entropy (CEE) to the free energy of a statistical-mechanics model, where noise and QE operation exactly correspond to two symmetry-breaking fields in different directions. Combining with unitary gate and measurement, we obtain a random field model with ferromagnetic coupling. This model exhibits a ferromagnetic-paramagnetic phase transition driven by both increasing temperature and increasing random field strength, under the condition of zero net external field. This phase transition thus leads to the transition in the scaling of CEE. Next, we discover that the zero-field condition is equivalent to an average apparatus-environment exchange (aAEE) symmetry, which ensures the CEE to be a valid probe of entanglement [43, 44]. Consequently, the transition observed in CEE directly corresponds to an entanglement phase transition. By imposing the aAEE symmetry, we conduct numerical studies for (2+1)-d circuits and explicitly demonstrate the MIPT in the presence of various decoherence noises. Our approach is inspired by Ref. [43], where a prototype setting was introduced. The structure we introduced here incorporates measurement and is much easier to implement in practice, as will be further elaborated later.
@@
Analytical analysis. Instead of considering subsystem entanglement entropy, we aim to calculate the CEE $\overline{S(M|A)}\equiv\overline{S(M,A)}-\overline{S(A)}$. $M$ denotes an arbitrary subsystem and the overline here represents averaging all circuits and measurement outcomes. Mathematically, this can be expressed as,
@@
$\mathcal{F}^{n,k}=-\text{log}\mathcal{Z}^{(n,k)}$ is identified as the free energy [46]. The underlying statistical-mechanics model for $\mathcal{Z}^{(n,k)}$ consists of spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$, where $\mathbb{I}$ and $\mathbb{C}$ correspond to two particular group element in $\mathbb{S}(Q)$ [46]. A critical observation is that, in both $\mathcal{Z}_{MA}$ and $\mathcal{Z}_{A}$, the environment qubits are subject to the same boundary condition $\mathbb{I}$, while the ancilla qubits are subject to $\mathbb{C}$. Consequently, we can interpret them as competing external fields applied at random positions within the bulk, with $\Lambda_{A}^{(n,k)}$ representing the bulk partition function. In other words, through QE operation and considering the conditional entropy, a symmetry-breaking field along the $\mathbb{C}$ direction is effectively introduced to counteract the influence of noise. Moreover, the unitary gates induce a ferromagnetic coupling, while the measurement rate acting as the temperature [46]. Thus, $\overline{S(M|A)}$ is directly related to the free energy difference of a random field model with ferromagnetic coupling, subject to different boundary conditions, as shown in Fig. 2(a).
In the absence of random fields, the model exhibits spin permutation symmetry [47]. However, any non-zero net external field explicitly breaks this symmetry. As a result, the bulk spins become aligned with either $\mathbb{I}$ or $\mathbb{C}$, leading to the formation of domain walls due to the imposed boundary conditions. The free energy cost associated with these domain walls scales proportionally with the length of the subsystem, resulting in a volume-law scaling of the CEE. A transition from volume-law to area-law scaling becomes possible only when the net external field vanishes, restoring the spin permutation symmetry between $\mathbb{I}$ and $\mathbb{C}$. This requires that the QE operation is symmetric to the noise, along with equal rates for them. Specifically, unless otherwise indicated, we assume the symmetric scenario where $U_{SE}=U_{SA}$, $\rho_{E}=\rho_{A}$ and $q_{n}=q_{e}=q/2$ in the following. Given prior work demonstrating phase transitions in random field Ising models and Potts models with zero-field conditon [48, 49, 50, 51, 52, 53, 54, 55, 56, 57], we anticipate that a phase transition would occur at finite $p$ and $q$ in quantum circuit. At high temperatures (large $p$), the system enters a paramagnetic phase where domain walls are absent, resulting in area-law scaling. Conversely, at low temperatures (small $p$), the symmetry is spontaneously broken and domain walls form along the boundary, yielding volume-law scaling.
@@
Average symmetry. Since $\overline{S(M|A)}$, as opposed to $\overline{S(M)}$ in traditional setting, undergoes a phase transition, it is pertinent to determine under what conditions $\overline{S(M|A)}$ serves as a valid probe of entanglement or quantum information [58]. This ensures that the observed phase transition constitutes a genuine entanglement phase transition with meaningful operational significance. It has been proven that if the system has information exchange (IE) symmetry in every realization, such that $S(M|A)=S(M|E)$, then $S(M|A)$ ensures entanglement between $M$ and $M^{c}\cup A$ [44]. Here, $M^{c}$ denotes the complement of the subsystem $M$. Volume law scaling of $S(M|A)$ would then yield an extensive number of distillable Bell pairs between $M$ and $M^{c}\cup A$. The IE symmetry can be guaranteed if an apparatus-environment exchange (AEE) symmetry is present, which requires that each circuit realization be symmetric under the exchange of ancilla and environment qubits. However, this symmetry is explicitly violated by any realization in our circuit since noise and QE operation are independently chosen and applied at each location. Additionally, the number of ancilla and environment qubits may differ within a specific circuit realization, making the “exchange” operation more subtle to define. To solve this issue, we propose that an aAEE symmetry can still lead to an average IE (aIE) symmetry, thereby making the averaged CEE a valid probe of entanglement. The relationship between these symmetries are summarized in Fig. 3. We first define a generalized SWAP operation $\text{SWAP}^{\prime}$ on $\left|\Psi_{m}\right\rangle$ as initially adding unentangled qubits to either $E$ or $A$ to equalize the number of qubits. Subsequently, a conventional SWAP gate is applied to $E$ and $A$, followed by the removal of the unentangled qubits. $\text{SWAP}^{\prime}\left|\Psi_{m}\right\rangle$ would not necessarily equal to $\left|\Psi_{m}\right\rangle$, but may transform it into another realization $\left|\Psi_{m^{\prime}}\right\rangle$. The aAEE symmetry requires that the probabilities of these two realizations occurring are identical. Concretely, one can represent the circuit’s outcome as $\rho=\sum_{m}\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|$ [59], and the aAEE symmetry necessitates that
@@
This is reminiscent of the recently proposed average (weak) symmetry in the study of mixed state order [60, 61, 62, 63, 64, 65, 66]. With aAEE symmetry and combining with Eq. (2), one would have
$\sum_{m}\Lambda_{A}^{Q}$ $=$ $\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\text{SWAP}^{\prime\dagger}\right.$ (4)
$\left.\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It also ensures that $\overline{S(M|A)}$ represents the entanglement between $M$ and $M^{c}\cup A$.
Translating the aAEE symmetry into the concrete circuit model reveals that it actually coincides with the zero-field condition, which is necessary for satisfying Eq. (3). Therefore, the MIPT captured by CEE indeed represents the transition in the entanglement scaling law of the underlying quantum state. Moreover, since any noise can be represented in the Stinespring form [45], QE operations are capable of protecting MIPT from various types of noise.
@@
An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting. Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area law phase, then sweep through different QE operation rates $q_{e}$. As long as the circuit is deep in the area law phase, the bipartite CEE should not change upon varying $p$, once the condition $q_{n}=q_{e}$ is met. Therefore, curves of CEE corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$ in the thermodynamic limit. By performing finite-size scaling on the intersection point, we obtain an estimate of the noise rate $q_{n}$ in the system. A more detailed explanation and finite size numerical results are provided in [46].
Besides noise, post-selection is another significant challenge preventing MIPT from being easily observed. Here the post-selection problem persists when the measurement rate is non-zero because determining the CEE of a single trajectory requires post-selecting on the measurement outcomes, a difficulty rooted in Born’s rule. Recent work has utilized various post-selection-free methods to probe the phase transition, such as using cross-entropy benchmarking [72, 73] or providing bounds on the entanglement entropy [74, 75]. We believe these methods can also be extended to the conditional entropy in this work, and we leave this exploration for future work.
@@ Acknowledgements.
Acknowledgment. We thank Shane P. Kelly for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the National Key Research Program of China under Grant No. 2019YFA0308404, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.